Meteor Infrasound A meteor that explodes n the atm
Answer #1

d = 8830 km = 8830,000 m ; t = 8.51 h = 31248 s

lambda = 6.6 km = 6600 m

we know that,

speed = dist/time

v = 8830000/30636 = 282.58 m/s

we know that,

v = f lambda

f = v/lambda = 282.58/6600m = 0.0428 Hz

Hence, f = 0.0428 Hz.